# Making Every Frame Matter: Continuous Video Understanding for Large Models via Adaptive State Modeling

### v1

@@ Abstract.
Video understanding has become increasingly important with the rise of multi-modality applications. Understanding continuous video poses considerable challenges due to the fast expansion of streaming video, which contains multi-scale and untrimmed events. We introduce a novel system, C-VUE, to overcome these issues through adaptive state modeling. C-VUE has three key designs. The first is a long-range history modeling technique that uses a video-aware approach to retain historical video information. The second is a spatial redundancy reduction technique, which enhances the efficiency of history modeling based on temporal relations. The third is a parallel training structure that incorporates the frame-weighted loss to understand multi-scale events in long videos. Our C-VUE offers high accuracy and efficiency. It runs at speeds ¿30 FPS on typical edge devices and outperforms all baselines in accuracy. Moreover, applying C-VUE to a video foundation model as a video encoder in our case study resulted in a 0.46-point enhancement (on a 5-point scale) on the in-distribution dataset, and an improvement ranging from 1.19% to 4% on zero-shot datasets.
1. Introduction
Video understanding (madan2024foundation, ; chen2024video, ; huang2018makes, ; zolfaghari2018eco, ; wu2021towards, ; tang2023video, ) is an important and long-standing application for edge and mobile computing. It requires capturing spatio-temporal features from an input video to conduct high-level tasks, such as action localization, activity recognition, video question answering, and video captioning. These tasks are widely used in edge scenarios, including robotics, augmented reality, embedded devices (e.g., Kinect and Wii remote), and IoT devices. Due to user data privacy, responsiveness requirements, and the high cost of video transmission to the cloud, real-world video understanding is typically processed on edge devices.
Challenges. Compared to video processing tasks only relying on frame spatial features (wu2020emo, ; tong2021large, ; jiang2021flexible, ; voigtlaender2019mots, ; yi2020eagleeye, ; wu2021pecam, ), e.g., object detection, segmentation, and tracking, video understanding on edge is much challenging due to the following reasons.
(i) Video understanding needs to capture the spatio-temporal features of an event on different temporal scales (zhao2018trn, ; shao2020finegym, ). Figure 2 shows an indoor activity clipping from Charades dataset (sigurdsson2016hollywood, ). The temporal durations of activities vary greatly from around 155 frames (putting a bag) to 553 frames (making a sandwich). The challenge lies in modeling this variability effectively: shorter events demand quick, precise attention, while longer activities require retaining relevant context over extended periods, complicating feature extraction across different timescales.
@@ 1. Introduction
(ii) Video understanding normally needs to capture meaningful features from the raw streaming videos, which are untrimmed and dominated by noise and redundancy in both spatial and temporal dimensions. As seen in Figure 2, the pixels that are related to an activity are very sparsely distributed in the video. The understanding may be distracted by the dominant unrelated pixels.
@@
(iii) Video understanding on edge must guarantee a timely understanding of the scene while managing the resource constraints of typical edge devices. For instance, robots or embodied AI (duan2022survey, ) systems must continuously perceive and comprehend dynamic scenarios through video input to carry out reasoning (strafforello2023current, ) and task planning (gadre2022continuous, ).
Figure 1. C-VUE: continuous video understanding on streaming video. As each frame is received, C-VUE updates its hidden state and outputs the events that have occurred so far.
Figure 2. We utilize the video stream, along with annotations from the Charades dataset (sigurdsson2016hollywood, ), to illustrate our motivation. C-VUE aims to understand ”What has happened so far?” at any moment in a streaming video.
@@
Figure 3. Typical designs to process frames of the video stream for continuous video understanding. $\texttt{E}_{i}$ signifies an event while $\texttt{f}_{i}$ depicts the frame at time $t_{i}$. The “All Frames at Once” design processes all video frames at every moment. The “Sliding Window” design incorporates only the latest $m$ frames. The “State Assistance” relies on a single frame at a time and archives previous data through a hidden state. Our adaptive state modeling is an improved state assistance design, excelling in long-range history modeling via a video-aware technique.
Current status of technology. Past works explore different DNN methods for video understanding, including 3D CNN (tran2015learning, ), RNN (e.g., LSTM) (hochreiter1997long, ), Transformer (vaswani2017attention, ), and the recent Mamba (gu2023mamba, ). However, none of the past works can mitigate all the challenges at the same time.
To capture the features of various temporal scales, a straightforward method is to input all the past frames into the model at each time step, as illustrated in Figure 3(a). However, it is certainly too costly for model inference and training. Take a Transformer model as an example, with the number of streaming frames $n$ grow, the computation cost of video understanding is $O(n^{3})$. Even with the many linear alternatives of Transformers, such as Linformer (wang2020linformer, ), RetNet (sun2023retentive, ), Mamba (gu2023mamba, ), and RWKV (peng2023rwkv, ), the cost is still $O(n^{2})$.
The most commonly used technique to control the cost is to fix the input size of the model even with the growing number of frames, such as sliding window or sampling, shown in Figure 3(b). For example, TSN (wang2016temporal, ) first divides the past video into segments and selects a random snippet from each segment for activity recognition. VideoLlama2 (damonlpsg2024videollama2, ) only samples 8 or 16 frames as input for any video length. This method has to ignore many detailed features.
There are also works (shi2015convolutional, ; huang2020efficient, ; zhou2018temporal, ), shown in Figure 3(c), that only input one current frame into a model (e.g., CNN) first to extract spatial features, and then conduct feature aggregation (e.g., through pooling (kar2017adascan, ) or LSTM (hochreiter1997long, )) among frames over time to capture temporal dependencies by hidden states. This is the most efficient method for inference and training, and all frames can be processed with no one lost.
This method, however, struggles to capture the temporal relations in long range, as the fundamental limitation discussed in (shao2020finegym, ; wang2016temporal, ). Current works’ limitations highlight the difficulty in creating a continuous video understanding method that meets the above three challenges. An ideal strategy should keep a fair balance between the accuracy of the Transformer and the efficiency of the LSTM, interpreting multi-scale events in rapidly growing videos. 1) It needs to correctly keep track of past state when forming decisions. 2) It must operate in real time on typical edge devices, importantly, without reducing processing speed as video length expands.
Key techniques. To address the challenges, we propose and implement C-VUE, a novel video encoder and the according training and inference system. C-VUE can perform real-time and prompt video understanding on streaming video in a continuous manner. That is to say, as each frame is received, the system updates its hidden state and outputs the events that have occurred so far, shown in Figure 1. Our C-VUE includes three key techniques. First, a long-range history modeling, designed to maintain the video historical context via the hidden state through a video-aware idea. Second, the temporal-relation guided spatial redundancy reduction, crafted to build a more efficient video modeling at the feature level. Third, a parallel training structure and a frame-weighted loss for the understanding of multi-scale events in the streaming video.
Evaluation results. We evaluate C-VUE across various downstream tasks, comparing it with widely-used models for video understanding, including Transformer (vaswani2017attention, ), RNN (LSTM) (hochreiter1997long, ), CNN (kar2017adascan, ), as well as the SOTA video understanding system for edge scenarios, i.e., MoviNetA0 (kondratyuk2021movinets, ) and Mvitv2 (li2022mvitv2, ). The evaluation is conducted to demonstrate five capabilities: (i) the understanding of events with various temporal lengths; (ii) prompt recognition of an event as more frames are received; (iii) generality to long-tail events; (iv) the localization and understanding of continuous input events; (v) long-range memorization.
The real-world video datasets used are FineGym (shao2020finegym, ) (sports gymnasium dataset), and Charades (sigurdsson2016hollywood, ) (daily indoor activities). We also compose more challenging datasets, e.g., concatenating all videos from the FineGym dataset into one to generate an hour-long video stream. Results show that with significantly fewer trainable parameters (ranging from 8.5k to 34.1M less than the baselines), C-VUE achieves accuracy improvements of 1.2% to 4.7% over Transformer, 2.6% to 5.9% over ConvPooling, 2% to 3% over LSTM, 14.5% to 67.4% over MoviNetA0, and 13.1% to 79.7% over MViTv2.
@@
Regarding latency, we evaluate C-VUE on typical edge devices (NVIDIA Jetson Orin, Apple MacBook, and an Intel PC with NVIDIA 4090 GPU). It achieves ¿30FPS on all these devices, up to 20$\times$ faster than the available inference systems.
@@
For the case study, we apply C-VUE video encoder to a video foundation model (Section 6). Specifically, we replace the visual encoder of the SOTA visual language model VideoLlama2 (damonlpsg2024videollama2, ) by C-VUE. The results demonstrate a 0.46-point improvement (on a 5-point scale) on the in-distribution dataset, and an improvement ranging from 1.19% to 4% on zero-shot datasets.
Contributions. We summarize our contribution as follows:
@@
  • We propose a continuous video understanding system, C-VUE, based on adaptive state modeling. It promptly understands incoming video frames by maintaining historical context using a hidden state.
  • C-VUE employs long-range history modeling to utilize the fixed-size hidden state for memorizing extensive historical data, reduces spatial redundancy based on temporal relations, and adopts a parallel training structure along with a frame-weighted loss function.
@@
  • Experiments show that C-VUE surpasses all baselines in continuous video understanding, using fewer parameters and achieving over 30 FPS on typical edge devices. A case study demonstrates that C-VUE can enhance the video foundation models on video QA tasks by serving as the video encoder.
2. Background
Continuous video understanding holds significant implications for complex human-computer interaction, embodied AI, and robots (duan2022survey, ; strafforello2023current, ; gadre2022continuous, ). While various approaches strive to achieve continuous understanding, none have solved this challenging task.
All Frames at Once. (Figure 3(a)) This method performs the video understanding on all video frames, theoretically yielding the highest task accuracy (slowfast2019fei, ; kondratyuk2021movinets, ). However, its efficiency is a concern. A basic transformer model has complexity $\mathcal{O}(n^{2})$, making the whole process have $\mathcal{O}(n^{3})$ complexity. Applying optimized architecture can decrease it to $\mathcal{O}(n^{2})$. Our experiment (Figure 4) shows that as frames increase, FPS drops to 0.05 times, a 95% decrease.
Fixed Window Size. (Figure 3(b)) This approach uses a fixed window size, like sliding window or sampling, to lower computational burden (Lin_2019_ICCV, ; Jiang_2019_ICCV, ; Li_2020_CVPR, ; li2022mvitv2, ). More efficient than the first method, its complexity ranges from $\mathcal{O}(mn)$ to $\mathcal{O}(m^{2}n)$, where $m$ equals the window size. The main challenge here is choosing an appropriate window size, as real-world events vary in duration. Small windows might miss longer events, and large ones could hinder understanding of shorter events.
State Assistance. (Figure 3(c)) By maintaining a group of hidden states, this method significantly enhances efficiency by reducing complexity to approximately $\mathcal{O}(n)$ (shi2015convolutional, ; huang2020efficient, ). However, as the input size grows, the fixed-size state struggles to remember the increasing history of inputs. Existing designs handle states in a video-unaware manner, preventing full utilization of the hidden states and the inability to model long-range history.
@@ 2. Background
Figure 4. Processing speed of understanding a 100k-frame video on RTX4090. Baselines 1(A) and 2(A) represent Transformer and Convpooling-based solutions using the “All Frames at Once” design. Baselines 1(F) and 2(F) employ the “Fixed window size” design (the window size is 100). Baseline 3 denotes an LSTM-based solution leveraging “State Assistance” design.
3. C-VUE Design
3.1. Design Overview
Design Requirements. To continuously process the video stream, an ideal design should integrate the advantages of the aforementioned approaches. The first advantage is efficiency: input only the current frame instead of repeating or tracing back through previous frames (similar to the ”State Assistance” design). The second advantage is accuracy: have a comprehensive memory of historical information at all times (approximating the ”All Frames at Once” design).
Design Challenges. To achieve the ideal design, the key lies in managing the growing volume of video, including: First, maximizing the memorization of the historical video data in a bounded subspace. Second, enabling processing at the edge in real-time without any increase in processing time as the video length extends.
Design Preliminary. We base our method on a state-assistance design. The typical state-assistance design involves recurrent updates over a sequence through hidden states (sieber2024understanding, ). We take the RNN as an example to introduce the standard state-assistance design. It computes $\mathbf{y}$ through a dynamic recurrence of input signals $\mathbf{x}$ at each time step $t$ through hidden states $\mathbf{s}$. Formally,
(1) $s_{t}=\phi_{h}(\mathbf{A}s_{t-1}+\mathbf{B}x_{t}+\mathbf{b}_{h}),$
(2) $y_{t}=\phi_{o}(\mathbf{C}s_{t}+\mathbf{b}_{o}),$
where $\mathbf{A},\mathbf{B}$, and $\mathbf{C}$ denote the state matrix, input matrix, and output matrix, respectively. The $\mathbf{b}_{h}$ and $\mathbf{b}_{o}$ are the bias parameters, and the functions $\phi_{h}$ and $\phi_{o}$ are activation functions.
Subsequently, the gating mechanism is introduced to enhance the memory capability of state-assistance designs, e.g., LSTM (sherstinsky2020fundamentals, ). More specifically, gates denoted as $g^{s}_{t}$, $g^{i}_{t}$, and $g^{o}_{t}$ are introduced at each state, input, and output, respectively. At time $t$, these gates dynamically adjust the state update process and the output based on the given input. Formally,
(3) $s_{t}=\phi_{h}(g^{s}_{t}\circ\mathbf{A}s_{t-1}+g^{i}_{t}\circ\mathbf{B}x_{t}+\mathbf{b}_{h}),$
(4) $y_{t}=\phi_{o}(g^{o}_{t}\circ\mathbf{C}s_{t}+\mathbf{b}_{o}).$
This gating mechanism resembles a weighting scheme, represented as a vector with values ranging from 0 to 1, which is computed based on the input and the last output.
Limitation of RNN and Its Variants. The RNN and its gating-based variants are widely considered difficult to train and challenging to handle long-range historical data by many research works (greaves2019statistical, ; zhao2020rnn, ; miller2018stable, ). Our experimental results also indicate that the LSTM-based baseline performs poorly on hours-long videos (Table 2).
We argue that the underperformance of such methods is due to their inefficient utilization of the hidden state. Specifically, standard RNN processes input at different moments identically, i.e., fixed $A$, $B$, and $C$ in Equations 1 and 2, resulting in a lack of adaptation towards these inputs. When redundancies exist within the input data, this leads to an irrational use of the hidden state. Despite the subsequent introduction of gating mechanisms to enhance input adaptation, these designs merely adopt a weighted approach to decide how the new inputs should merge with the existing states. Such a design idea lacks flexibility and fails to fully tap into the memorization capability of the hidden state.
Our Key Designs. We introduce our C-VUE with the following three parts. First, long-range history modeling (Section 3.2). This is our core design, aiming to maximize the memorization capability of multi-scale events over long sequences. This design adopts two techniques: 1) a fully video-aware mechanism to improve the hidden state’s utilization; and 2) a scale invariance mechanism for multi-scale events understanding.
Second, temporal relation-guided spatial redundancy reduction (Section 3.3). Performing history modeling in pixel space can result in excessive system overhead and increased learning difficulty. This design enables the efficient modeling by eliminating the redundant information at the feature level with the guidance of the temporal relation.
Third, parallal training with frame-weighted loss (Section 3.4). It introduces how our C-VUE builds the training structure for continuous long video understanding. We detail the efficient parallel training for long video samples and the weighted loss design for the understanding of every frame.
3.2. Long-range History Modeling
The challenge of understanding long-range history primarily stems from two aspects, namely inter-event noise and intra-event sparsity. The inter-event noise implies that there may be numerous noise frames between two events. In the context of continuous understanding, the longer the video, the more pronounced this issue becomes. The intra-event sparsity represents the inherent feature sparsity of event information. Some events carry limited information despite their extended durations, e.g., sleeping. However, the multi-scale characteristics of events make it difficult to handle this sparsity.
To handle these issues, we propose two techniques. The first is a fully video-aware history modeling. Here, each fixed parameter is improved to a dynamically-generated parameter based on the input, enhancing the ability to filter noise between events. The second is a scale-invariant event memorization technique, which dynamically determines how an event should be memorized in the hidden state based on its information density, rather than its duration.
3.2.1. Fully Video-aware History Modeling
In existing designs, once trained, the model parameters cannot be flexibly adjusted based on the input frame. Intuitively, if some frame is repetitive or does not contain any events, the design should discard this frame without updating the hidden state, and simply copy the result from the last output. Therefore, we aim to improve all matrix parameters into input-dependent ones. Our design incorporates dynamic adaptability, similar to the attention mechanism in transformers. The attention mechanism adapts to different inputs by adjusting its focus on various parts of the input sequence.
Many efforts have explored that, to model the long-range history, the matrix $A$ should be designed carefully. So, in our design, we only improve matrics $B$ and $C$ to input-dependent ones. Then we detail how to improve the matrix $A$ for long-range history modeling.
We improve the $B$ and $C$ as functions of the input, i.e., $B_{t}=f_{B}(x_{t})$ and $C_{t}=f_{C}(x_{t})$. Consequently, during training, we no longer learn the $B$ and $C$ matrices directly but instead learn how to dynamically generate these matrices based on the input. The hidden state is maintained dependent on the input $x$ to only keep the event features. In our method, we utilize the two linear layers as the function $f_{B}$ and the function $f_{C}$, respectively.
Matrix $A$ controls how information from previous states is transferred to the new state. The design of matrix $A$ determines whether the method retains only recent states (local dependency) or captures all past states. Recently, many efforts have been put into how to improve the ability to memorize long-range history by designing matrix $A$.
Inspired by Work (gu2023mamba, ), we redesign the transform matrix $A$ and its initialization method. First, we enhance matrix $A$ by making it a diagonal matrix, which has been proven to significantly improve the capability to model long-range history. Various efforts have explored alternative initialization techniques for matrix $A$ during training to boost model performance. Following this work, we employ random initialization techniques, which achieve similar results. Therefore, in our C-VUE, we utilize a randomly initialized diagonal matrix $A$ to enable long-range memorization.
3.2.2. Scale-invariance Events Memorization
As shown in Figure 2, different events contain varying numbers of frames, but they may convey a similar amount of information. We aim to use hidden states of similar sizes to memorize events with the same amount of information, even though these events have different time scales.
Inspired by Work (nguyen2022s4nd, ), we can leverage discretization techniques to dynamically resample video frames. The input frames $x_{0}$, $x_{1}$, …, $x_{t}$ can be viewed as uniformly sampled from an underlying signal $f(x)$ with a step size $\Delta$, i.e., $x_{t}=f(t\Delta)$. The standard RNN utilizes matrics $A$ and $B$ to update the state in keeping with $\Delta$ (Equation 1 and 2).
To achieve our design goal, we no longer use video frames as the basic unit for sampling but instead use the events as the basic unit for sampling via a dynamic $\Delta$. Then, we apply the dynamic $\Delta$ on the state update procedure to achieve the dynamic resampling. Note that the step size $\Delta$ is not a real sampling rate, and we do not perform the resampling technique on the video frames. All input frames are processed during the continuous video understanding procedure.
We utilize the discretization technique, i.e., zero-order hold (ZOH), utilized in (gu2023mamba, ) to achieve the above dynamic resampling. Specifically, $A$ and $B_{t}$ can be improved with an input-dependent parameter $\Delta_{t}=f_{\Delta}(x_{t})$:
(5) $\overline{A_{t}}=\exp(\Delta_{t}A)$
(6) $\overline{B_{t}}=(\Delta_{t}A)^{-1}\exp(\Delta_{t}A-I)\cdot\Delta_{t}B_{t}$
$\Delta_{t}$ is calculated based on the input $x_{t}$, thus achieves the awareness of the current event. We use a linear layer as the $f_{\Delta}$ function.
3.3. Temporal Relation-guided Spatial Redundancy Reduction
Videos contain both spatial (within-frame) and temporal (between-frame) information, both critical for accurate understanding. However, as shown in Figure 5, video data contains a vast amount of redundant information in spatial space. Directly performing history modeling introduces lots of redundant computations, leading to poor system performance.
Eliminating the redundancy is challenging. Because, firstly, the event feature is sparse. The event feature in a video sample may be captured by only a small portion of the pixel content. Secondly, object movement causes poor locality at the pixel level. It is hard to accurately understand events merely by analyzing two adjacent regions (by setting thresholds). Therefore, if we perform video understanding at the pixel level like traditional approaches, e.g., MoviNet (kondratyuk2021movinets, ) and Mvitv2 (li2022mvitv2, ), there are lots of redundant computation (as shown in the left part of Figure 6).
@@ 3.3. Temporal Relation-guided Spatial Redundancy Reduction
Figure 5. Video frames of the balance beam sports, selected from the FineGym99 dataset (shao2020finegym, ). The entire gymnastic movement was captured in only 17 frames, lasting less than 700 milliseconds.
Figure 6. Different ideas to model the video information. Traditional methods associate any two pixels or patches with complexity $\mathcal{O}(n^{2})$ or $\mathcal{O}(m^{2})$, where $n$ and $m$ are the pixel and patch counts. By reducing spatial redundancy, we model video content more efficiently.
Given the issues mentioned earlier, many pixels, such as background pixels in Figure 5, do not contain event information, leading to computational redundancy. Therefore, such a method inevitably leads to lots of computation redundancy. Additionally, due to low inter-frame pixel locality caused by object movement, it is challenging to reduce the quadratic complexity simply by searching nearby pixels.
Traditional video processing methods have carried over into the era of LLM-based video foundation models (damonlpsg2024videollama2, ). The vision encoder break a frame into patches and then correlate these patches across frames in an all-to-all manner. The patch number is $m$, and the complexity is $\mathcal{O}(m^{2})$. This approach makes it difficult for video foundation models to balance accuracy and system performance. Correlating a large number of frames is costly in both the training and inference phases. Specifically, the high computational complexity often requires excessive downsampling, which reduces accuracy. We analyze this issue in Section 6 and show how our design can improve video foundation models.
Our approach introduces a temporal relation-guided spatial redundancy reduction. The key insight is that we can learn the rule to reduce the redundant spatial information, that is not related to the event, by correlating these features in the temporal dimension. For example, understanding a jumping action, as shown in Figure 5, is primarily about observing how the person’s pose and position change between frames, which is unrelated to what the person is wearing. Hence, we can first attempt to capture the specific pose and position information of the objects at each moment, then associate them over time for further understanding.

### v2

# Making Every Frame Matter: Continuous Activity Recognition in Streaming Video via Adaptive Video Context Modeling

@@ Abstract
Video activity recognition has become increasingly important in robots and embodied AI. Recognizing continuous video activities poses considerable challenges due to the fast expansion of streaming video, which contains multi-scale and untrimmed activities. We introduce a novel system, CARS, to overcome these issues through adaptive video context modeling. Adaptive video context modeling refers to selectively maintaining activity-related features in temporal and spatial dimensions. CARS has two key designs. The first is an activity spatial feature extraction by eliminating irrelevant visual features while maintaining recognition accuracy. The second is an activity-aware state update introducing dynamic adaptability to better preserve the video context for multi-scale activity recognition. Our CARS runs at speeds $>$30 FPS on typical edge devices and outperforms all baselines by 1.2% to 79.7% in accuracy. Moreover, we explore applying CARS to a large video model as a video encoder. Experimental results show that our CARS can result in a 0.46-point enhancement (on a 5-point scale) on the in-distribution video activity dataset, and an improvement ranging from 1.19% to 4% on zero-shot video activity datasets.
1 Introduction
Video activity recognition [26, 7, 16, 48, 42, 37] is an important and long-standing application. These tasks are widely used in advanced human-computer interactions (HCI), including robotics and augmented reality [9, 35, 11]. To enable advanced HCI, real-time recognition of streaming video on resource-constrained mobile or edge devices is crucial. Activity recognition relies heavily on video context, making it difficult to determine actions from a single frame. Therefore, a continuous video activity recognition system must maintain video context, allowing it to infer activities in new frames based on the accumulated context. However, achieving efficient continuous recognition in streaming video is challenging for the following reasons.
Figure 1: An example of continuous video activity recognition in streaming video. CARS takes every frame as input and continuously recognizes human activities.
(i) Diverse Temporal Scales of Activities. Continuous video activity recognition needs to capture the spatio-temporal features of activity on multiple temporal scales [46, 30]. As shown in Figure 1, activities at different scales may have dependencies or causal relationships, such as taking bread from a bag as part of making a sandwich. The system must continuously recognize activities of varying lengths based on newly received video frames.
@@ 1 Introduction
(ii) Accumulating Redundancy in Streaming Video. Continuous video activity recognition normally needs to capture meaningful features from the raw streaming videos, which are dominated by noise and redundancy. As illustrated in Figure 5, activity-related pixels are sparsely distributed in the video. Without effective online reduction, the accumulation of redundant features can quickly overwhelm the activity-related features as the video progresses.
@@
(iii) Computational Efficiency Under Resource Constraints. Continuous video activity recognition must guarantee a timely recognition of the scene while managing the resource constraints of typical edge devices. For instance, robots or embodied AI [9] systems must continuously perceive dynamic scenarios through streaming video input to carry out reasoning [35] and task planning [11].
Existing video recognition efforts have explored numerous strategies for maintaining video context, yet none have been able to address all the challenges simultaneously. Works, using “all frames at once” strategy (Figure 3(a)), perform activity recognition on every frame of the video, which theoretically yields the highest accuracy [10, 19]. However, its efficiency is the biggest concern. Other works use “sliding window” strategy (Figure 3(b)) to reduce computational cost [23, 17, 21, 22]. However, these methods struggle to accurately recognize activities at different scales. Some approaches use a “hidden state assistance” design (Figure 3(c)), where a hidden state compresses video context to enable more efficient activity recognition. However, existing state-assisted methods do not effectively eliminate spatiotemporal redundancy in the streaming video, resulting in inefficient utilization of the hidden state and low accuracy.
Figure 2: Difference between our idea and existing solutions on maintaining the hidden state.
The all-frame-at-once approach suffers from efficiency degradation over time as data volume grows, while the sliding window method often faces challenges in selecting optimal window sizes or sampling strategies, compromising accuracy. Therefore, our design builds on the existing state-assistant framework, which is efficient—processing one frame at a time—and accurate—adaptively maintaining video context through carefully designed hidden states. As illustrated in Figure 2, we argue that existing state-assistant works lack activity-aware design in both temporal and spatial dimensions. They tend to compress all video features, including the redundant background features, into the hidden state, resulting in inefficient use of hidden states, poor preservation of video context, and low recognition accuracy. Our core idea is to selectively compress only activity-related features in temporal and spatial dimensions, which we name adaptive video context modeling.
@@
Figure 3: Typical designs for streaming video activity recognition. The “All Frames at Once” design processes all video frames at every moment. The “Sliding Window” design incorporates only the latest (or sampled) $m$ frames. The “State Assistance” relies on a single frame at a time and maintains video context through a hidden state.
In this paper, we propose CARS for continuous video activity recognition in the streaming video, utilizing the adaptive video context modeling, with two key designs. First, activity spatial feature extraction (Section 3.1). To reduce redundant visual information, we employ a data reduction strategy that operates without requiring supervision from masks or bounding boxes of activity-related features. During training, we maximize feature elimination from the video input while preserving recognition accuracy. This adversarial learning process allows our CARS to discard spatial features irrelevant to the activity. Second, activity-aware state update (Section 3.2). Existing methods struggle to effectively maintain video context for multi-scale activities. Our design integrates a dynamic adaptability, akin to the attention mechanism in transformers. We dynamically recalculate the weight parameters based on the current input frame during the input processing, hidden state update, and output generation.
@@
We evaluate CARS across various activity recognition tasks, comparing it with widely-used models for video recognition. Results show that with significantly fewer trainable parameters (ranging from 8.5k to 34.1M less than the baselines), CARS achieves accuracy improvements of 1.2% to 79.7% over baseline methods. Regarding latency, we evaluate CARS on typical edge devices (NVIDIA Jetson Orin, Apple MacBook, and an Intel PC with NVIDIA 4090 GPU). It achieves $>$30FPS on all these devices, up to 20$\times$ faster than the available inference systems.
@@
To further explore the capability of our design, we apply CARS as a video encoder to a large video model to improve its activity recognition capability (Section 5). Specifically, we replace the visual encoder of the SOTA large video model VideoLlama2 [8] by CARS. The results demonstrate a 0.46-point improvement (on a 5-point scale) on the in-distribution dataset, and an improvement ranging from 1.19% to 4% on zero-shot datasets.
Contribution. We summarize our contribution as follows:
@@
  • We propose a continuous video activity recognition system, CARS, based on adaptive video context modeling. It promptly recognizes the activities in incoming video frames by efficiently maintaining video context for multi-scale activities.
  • Our CARS incorporates two key techniques to enhance the utilization efficiency of hidden states, i.e., minimizing redundancy from activity-irrelevant spatial information, and improving adaptability to streaming input while maintaining state.
@@
  • Experiments show that CARS surpasses all baselines in streaming video activity recognition, using fewer parameters and achieving over 30 FPS on typical edge devices, which can enable on-device video recognition for embodied AI. We also showcase that CARS can enhance the large video models by serving as the video encoder.
2 Background & Related Works
Various approaches strive to achieve continuous activity recognition, but none have solved this challenging task.
@@ 2 Background & Related Works
Figure 4: Processing speed of video activity recognition on RTX4090 (a 100k-frame video). Baselines 1(A) and 2(A) represent Transformer and Convpooling-based solutions using the “All Frames at Once” design. Baselines 1(F) and 2(F) employ the “Fixed window size” design (the window size is 100). Baseline 3 denotes an LSTM-based “State Assistance” design.
All Frames at Once. A straightforward method is to input all the past frames into the model at each time step, as illustrated in Figure 3(a). However, it is certainly too costly for model inference and training. Take a Transformer model as an example, with the number of streaming frames $n$ grow, the computation cost of video activity recognition is $O(n^{3})$. Even with the many linear alternatives of Transformers, such as Linformer [41], RetNet [36], Mamba [13], and RWKV [28], the cost is still $O(n^{2})$. Our experiment (Figure 4) shows that as frames increase, FPS drops to 0.05 times, a 95% decrease.
Fixed Window Size. The most commonly used technique to control the cost is to fix the input size of the model even with the growing number of frames, such as sliding window or sampling [23, 17, 21, 22], shown in Figure 3(b). For example, TSN [40] first divides the past video into segments and selects a random snippet from each segment for activity recognition. VideoLlama2 [8] only samples 8 or 16 frames as input for any video length. This method has to ignore many detailed features. VideoMamba [20] adopt the Mamba [13] as the architecture to perform the efficient video understanding, however, it still adheres to the batch-based processing strategy. The main challenge is selecting an optimal window size to balance the detection of both short and long activities.
State Assistance. There are also works [32, 15, 47], shown in Figure 3(c), that only input one current frame into a model (e.g., CNN) first to extract spatial features, and then conduct feature aggregation (e.g., through pooling [18] or LSTM [14]) among frames over time to capture temporal dependencies by hidden states. By maintaining a group of hidden states, this method significantly enhances efficiency by reducing complexity to approximately $\mathcal{O}(n)$ [32, 15]. However, as the input size grows, the fixed-size state struggles to remember the increasing history of inputs [30, 40].
@@
Figure 5: Video frames of the balance beam sports, selected from the FineGym99 dataset [30]. The entire gymnastic movement was captured in only 17 frames, lasting less than 700 milliseconds.
3 CARS Design
A motivating video sample. Video streams contain both spatial (within-frame) and temporal (between-frame) information, both critical for accurate recognition. Figure 5 presents a video sample from a real-world gymnastics match to motivate our design and explain the challenges we faced. There are two issues we can observe from the video sample. Issue 1: the activity-related spatial feature is sparse. The activity feature in a video sample may be captured by only a small portion of the pixel content. Issue 2: object movement causes poor locality at the pixel level. It is hard to accurately recognize activities merely by analyzing two adjacent regions (by setting thresholds).
We have observed the efficiency advantages of state-assistant methods in handling streaming video. Therefore, we utilize a state-assistant design, which involves recurrent updates over a sequence through hidden states [33]. Our CARS has two key designs, i.e., activity spatial feature extraction and activity-aware state update, to solve these two issues. Figure 6 illustrates the system workflow.
During the usage, streaming video is continuously input into the system frame by frame (without sampling). Our CARS extracts activity-related features and sends them to the activity-aware state update module. During the state update, the system attempts to identify possible activities and uses the activity features to update the corresponding part of the hidden state. If no further activity features related to those features maintained in the hidden state arrive, the corresponding video context is cleared from the hidden state. The updated state is stored in the system, waiting for the next frame to update it. Simultaneously, the updated hidden state can be sent to a classifier for timely activity recognition. With our optimization, the entire system can perform real-time video activity recognition on typical edge devices.
Figure 6: CARS’s workflow for continuous activity recognition via adaptive video context modeling by selectively compressing only activity-related features into the hidden state.
3.1 Activity Spatial Feature Extraction